State if the three side lengths form an acute, obtuse, or right triangle. ( 5 in, 12 in, 13 in ) *
1
Expert's answer
2020-05-22T15:03:02-0400

Solution.

"a=5 in;"

"b=12in;"

"c=13in;"

By the cosine theorem we find the angle against the larger side:


"с^2=a^2+b^2-2abcos\gamma;\implies cos\gamma=\dfrac{a^2+b^2-c^2}{2ab};"

"cos\gamma=\dfrac{25+144-169}{2\sdot5\sdot12}=0;"

"\gamma=90^o;"

Answer: The sides form a right angle.